OpenStudy (anonymous):

One of the roots of equation 2000x^6 +100x^5 +10x^3 +x -2 =0 is of form \[\frac{m+\sqrt{n}}{r}\] where m is non zero integer n and r are relatively prime natural nos.FInd value of m+n+r

OpenStudy (anonymous):

Don't Cheat

OpenStudy (anonymous):

Donot wolfram it

OpenStudy (anonymous):

I won't :-)

OpenStudy (anonymous):

\[\frac{m-\sqrt{n}}{r}\] must be root

OpenStudy (anonymous):

I got it to \[(100x^4 +10x^2+1)(20x^2-2 +x)=0\]

OpenStudy (anonymous):

Now I think we can solve it

OpenStudy (anonymous):

200 :P

OpenStudy (anonymous):

took me time to factorize :-/

OpenStudy (anonymous):

Yes, cinar is right I got 200 as well

OpenStudy (anonymous):

I love solving such questions, these are so fun. I hate school math

OpenStudy (anonymous):

\[(ax^2+bx+c)*(dx^4+ex^3+fx^2+gx+h)=\] =\[2000x^6+100x^3+x-2\]

OpenStudy (anonymous):

Hmm factorization that's all, take some terms common and then try to take some more that's it

OpenStudy (anonymous):

I like your way though

OpenStudy (anonymous):

wow then, you saw it huh..

OpenStudy (anonymous):

Did you use hit and trial. or some standard manipulatoin

OpenStudy (anonymous):

cinar's way is a standard way. mine is kind of hit and trial with inuition maybe

OpenStudy (anonymous):

Fact about polynomials is that any polynomial can be factorised into quadratic and linear factors but it is not always feasible

OpenStudy (anonymous):

Solivng a system of six equation doesnot seem promising also equatoins are not all linear

OpenStudy (anonymous):

hmm yeah. so what do you suggest what should we use?

OpenStudy (anonymous):

Cinar..we would find three sets of a,b,c But how to solve the system

OpenStudy (anonymous):

sorry guys, I have no time now, I can see it at night..

OpenStudy (anonymous):

Alright Meet me tommorrow 12:00 pm IST

OpenStudy (anonymous):

For more such problems Ishaan

OpenStudy (anonymous):

IST? Indian Standard Time? you from India?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

Yeah I like these but it's late here and I have a problem to finish. I have posted it on OpenStudy you can check it out http://openstudy.com/users/ishaan94#/updates/4f2d8011e4b0571e9cba7c3b

OpenStudy (anonymous):

Cool!